OpenStudy (anonymous):

What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thanks

OpenStudy (igreen):

Just solve for 's'. -7 + s + -14 = 12 - 7 - 14 Simplify: -7 + s + -14 = 12 - 21 Simplify: -7 + s + -14 = -9 Simplify: -21 + s = -9 Add 21 to both sides, what do you get?

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

my fault ....please look over igreen explanation to understand proplem